When a research balloon is released at sea level, where the temperature is 14.8°C

14.8°C and the atmospheric pressure is 1.01×10

5

 Pa

1.01×105 Pa, the helium in it has a volume of 12.9 m

3

12.9 m3. The balloon reaches an altitude of 32 km

32 km, where the temperature is −44.1°C

−44.1°C and the pressure is 867 Pa

867 Pa. What is the volume of the helium at this altitude? 


1
Expert's answer
2022-03-08T23:58:02-0500

T1=14.8+273= 287.8K

P1=1.01×105 Pa

V1=12.9 m3

T2=-44.1+273= 228.9K

P2=867Pa

V2=?

P1V1÷T1 = P2V2÷T2

1.01×105×12.9÷287.8 = 867×V2÷228.9

V2= 1.01×105×12.9×228.9÷(287.8×867)

=1195.22m3